OpenStudy (anonymous):

Find the critical numbers of the function. (Enter your answers as a comma-separated list. Use n to denote any arbitrary integer values. If an answer does not exist, enter DNE.) g(θ) = 36θ − 9 tan θ

OpenStudy (xapproachesinfinity):

hey @FATIMA955 so any idea where to start?

OpenStudy (xapproachesinfinity):

Do you know how to get critical numbers and values?

OpenStudy (xapproachesinfinity):

??

OpenStudy (xapproachesinfinity):

well find the derivative of \(g(\theta)\)

OpenStudy (xapproachesinfinity):

do it now!! hehe

OpenStudy (xapproachesinfinity):

Here something that you should know by now to find crictical point you need to find the derivative then set that derivative equal to zero and solve

OpenStudy (xapproachesinfinity):

Did you learn that in the lecture?

OpenStudy (xapproachesinfinity):

f is continuous on the interval [a,b] and differentiabale on (a,b) we say that c is a critical number if and only if f'(c)=0 or f'(c) does not exist

OpenStudy (xapproachesinfinity):

come on! do something please lol I can't help you if you are just watching the screen there

OpenStudy (xapproachesinfinity):

hey @FATIMA955, by now, you should be able to do any derivatives

OpenStudy (anonymous):

hey yeah sorry

OpenStudy (anonymous):

how am i suppose to know all derivates havent studied math for past 3 years yeah im done with the derivative part though looking at the tips u have given

OpenStudy (anonymous):

i need to find the value for theta thts it right?

OpenStudy (xapproachesinfinity):

oh ok, but if you are taking calc now you should review all those stuff again so \(\huge \rm g'(\theta)=-9sec^2(\theta) \) set it equal zero \(\huge \rm g'(\theta)=0 \Longrightarrow -9sec^2(\theta)=0\)

OpenStudy (xapproachesinfinity):

yes you need to solve for theta

OpenStudy (xapproachesinfinity):

you need some precalc here! trigonometry

OpenStudy (xapproachesinfinity):

so if we have \(\huge \rm -9sec^2(\theta)=0\) then \(\huge \rm sec(\theta)=0\) now you should ask when does sec give zero

OpenStudy (xapproachesinfinity):

you might need to relook how the graph of sec looks like

OpenStudy (anonymous):

this cant exist right

OpenStudy (xapproachesinfinity):

http://www.calculatorsoup.com/images/trig_plots/graph_sec_pi.gif

OpenStudy (xapproachesinfinity):

yes! secant can never be zero

OpenStudy (anonymous):

since cos x =1/sec x having it =0 would mean that the cosine is undefined for some value .so we conclude the ans doesnt exist??

OpenStudy (xapproachesinfinity):

So what do you think

OpenStudy (xapproachesinfinity):

hmm no! what we do is consider the values where g' doesn't exist meaning the values that make secant undefined it is not cosine who is undefined for some values but secant cosine is perfectly defined, you need to think what values make 1/cosx not defined

OpenStudy (anonymous):

after finding the dvt we get [36-9\sec^2\]

OpenStudy (xapproachesinfinity):

recall that f'(c) does not exist implies that c is critical point when you have f'(c)=0 c could be either maximum or minimum

OpenStudy (xapproachesinfinity):

no? i give you the derivative already

OpenStudy (anonymous):

yeah u did but it was 36 theta-9tan theta

OpenStudy (xapproachesinfinity):

No it was \(g'(\theta)=-9sec^2\theta\)

OpenStudy (xapproachesinfinity):

Are focusing here or not! i wrote it above that part is done! now we are looking for the points that make g' does not exist

OpenStudy (xapproachesinfinity):

the first thing i did is find \(\huge \rm g'(\theta)\) then i set up \(\huge \rm g'(\theta)=0\)

OpenStudy (xapproachesinfinity):

now we are in the final step to get the values of theta that make \(\huge \rm g'(\theta) ~undefined \)

OpenStudy (anonymous):

yes yes i get ittt..

OpenStudy (xapproachesinfinity):

if you look back here in the graph you should be able to tell what values make secant undefined which make g' also undefined

OpenStudy (xapproachesinfinity):

http://www.calculatorsoup.com/images/trig_plots/graph_sec_pi.gif

OpenStudy (xapproachesinfinity):

to make it easy for you the values are \(\huge \rm ...-5\pi/2-3\pi/2,-\pi/2,\pi/2,3\pi/2,5\pi/2...\)

OpenStudy (xapproachesinfinity):

the values continue from left and to right as well all of those values make secant undefined

OpenStudy (anonymous):

ohh ok thanksss..

OpenStudy (xapproachesinfinity):

well you are not finished yet lol and np!

OpenStudy (anonymous):

hahah yess i know ..

OpenStudy (xapproachesinfinity):

ok! good then

OpenStudy (xapproachesinfinity):

good luck i have to go

OpenStudy (anonymous):

oh ok sure thankss alot...

OpenStudy (xapproachesinfinity):

^_^